### v1

# technical note on calibrating vision-language models under covariate shift

Abstract
Despite being a successful example of emerging capability, vision-language foundation models for low-shot vision classification have a limited ability to sufficiently generalize to the target data distribution due to sample poverty, leading to sensitivity to variations in the data. A popular mitigation strategy is finetuning over multiple datasets, but domain generalization is expensive when practiced in this manner. This work examines both covariate shift between pre-training data and the underspecified target data, and confidence misalignment, where the model’s prediction confidence amplified by the limited data availability. We propose Confidence-Calibrated Covariate Shift Correction ($C3SC$), a unified framework to mitigate both covariate shift and confidence misalignment. $C3SC$ leverages Fisher information penalty for covariate shift correction and confidence misalignment penalty (CMP) to lower confidence on misclassified examples. Experimental results across various vision and covariate shift datasets demonstrates that $C3SC$ significantly improves in calibration (ECE) by $5.82\%$ at maximum. $C3SC$ shows better robustness as well by showing $3.5\%$ improvement in accuracy metric on challenging covariate shift datasets, making $C3SC$ a promising solution for reliable real-world vision-language low-shot applications under distribution shift.
1 Introduction
Foundation models like CLIP Radford et al. (2021), Flamingo Alayrac et al. (2022), Align Abdul Samadh et al. (2024) serve as the backbone of modern learning-based decision systems particularly for downstream vision classification tasks. However, real-world applications relying on these systems face two challenges. The first is, covariate shift, where the features distribution $P(x)$ changes between source and target data i.e $P(x)_{tr}\neq P(x)_{ts}$ while the conditional distribution remains unchanged $P(y|x)_{tr}=P(y|x)_{ts}$ Shimodaira (2000); Sugiyama and Müller (2005); Cortes and Mohri (2014); Khan et al. . This is being extensively studied under the domain generalization setting with foundation models Xiao et al. (2024); Murugesan et al. (2025); Wen et al. . The second is confidence misalignment, where the model,s $f(x)$ predicted confidence score does not reflect the likelihood of data Gal and Ghahramani (2016); Guo et al. (2017); Kumar et al. (2019). These issues are particularly prevalent in low-shot learning regimes, where limited labeled data lead to unstable target-dataset distributions. The distribution shift is inherent to the setting, and the low availability of target data unreliable predictionsKhan and Syed (2025) without that lack of certainty being reflected in predictive confidence.
Existing methods Khan et al. ; Murugesan et al. (2025); Wang et al. (2024); Khan and Syed (2025), address these problems individually. In this work, we present a unified solution which we refer Confidence-Calibrated Covariate Shift Correction ($C3SC$). Our framework incorporate Fisher information-based penalty the for correction of covariate shift and confidence misalignment penalty (CMP) to calibrate the confidence score. By integrating these penalties into contrastive loss of CLIP, $C3SC$ enables robust and reliable learning-systems in low-shot vision classification tasks.
Contributions
We investigate two different penalties offered by the literature for two different problems, and show that they could be convexly combined. The result shows that this augmented loss term addresses domain generalization while improving calibration, within a tight tolerance.
2 Method
The CLIP $\mathcal{L}_{\text{c}}$ uses contrastive loss for image-text align can be defined as: $\mathcal{L}_{\text{c}}=\frac{1}{2}\left(\mathcal{L}_{\text{txt}}+\mathcal{L}_{\text{img}}\right),$ while, text loss is given as: $\mathcal{L}_{\text{txt}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\left(\text{sim}(\mathbf{t}_{i},\mathbf{i}_{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\mathbf{t}_{i},\mathbf{i}_{j})/\tau\right)},$ and image loss is denoted as: $\mathcal{L}_{\text{img}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\left(\text{sim}(\mathbf{i}_{i},\mathbf{t}_{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\mathbf{i}_{i},\mathbf{t}_{j})/\tau\right)}$.
Covariate shift: In low-shot learning regimes, availability of limited data may differ training and test distributions, which leads to poor generalization. Simiarly, when data is distributed across batches or folds, each subset represents an extremely low-data setting, where distribution shift emerge due to insufficient sample representation. The information loss due to covariate shift caused by such fragmented data partitions Sugiyama et al. (2007); Moreno-Torres et al. (2012) can be effectively measured by Fisher information, just as in low-shot learning scenarios where model uncertainty is high due to limited observation Khan et al. ; vig (2010); Courtade (2016). The Fisher information measure can be defined as:
$I(\theta)=-\mathbb{E}\left[\frac{\partial^{2}\log P(X;\theta)}{\partial\theta^{2}}\right]$ (1)
Confidence misalginment: Foundation models such as CLIP often overconfident in predictions when image-text pair does not reconcile, specifically in low-shot settings Tu et al. (2024); Pandey et al. (2024); Murugesan et al. (2025); Khan and Syed (2025). CLIP contrastive loss maximizes $\mathcal{L}_{c}$ for correct image-text pair while minimizes it for incorrect pairs. Maximum entropy principle (MEP) defines that more uniform distributions best explain observed data Jaynes (1957). This principle contradicts with contrastive loss of CLIP, as $\mathcal{L}_{c}$ minimizes the loss for incorrect pairs $j\neq i$, the probability tends to zero which results in $H(P(\mathbf{t}|\mathbf{i};\theta))\rightarrow 0$ which results in highly peaked distribution. To overcome this problem confidence misalignment penalty (CMP) is introduce by Khan and Syed (2025) that is:
$\text{CMP}=\frac{P(x,y^{\prime})}{\sum_{y_{j}\neq y^{\prime}\,:\,P(x,y_{j})>P(x,y^{\prime})}P(x,y_{j})},$ (2)
where, $P(x,y^{\prime})$ is softmax probability for correct pairs and $P(x,y_{j})$ is for incorrect one.
CMP penalizes overconfident incorrect predictions by shifting log-likelihood to the true class.
Our method $C3SC$ integrates Fisher information penalty $I(\theta)$ and CMP into CLIP contrastive loss, which simultaneously correct covariate shift and ensure confidence calibration as well in foundation models.
Proposition 2.1.
The Fisher information $I(\theta)$ improves generalization by controlling the curvature of the loss landscape and CMP improves calibration by penalizing overconfidence and redistributing log-likelihood to the true class.
Proof.
The Fisher information $I(\theta)$ in eq 1 measures the curvature of log-likelihood function with respect to model parameters $\theta$. A high $I(\theta)$ indicates sharp curvature which results in poor generalization. To mitigate this effect, we introduce Fisher information regularization by adding $\lambda_{1}I(\theta)$ into the loss function: $\mathcal{L}_{\text{C3SC}}=\mathcal{L}_{\text{c}}+\lambda_{1}I(\theta).$ From PAC-Bayes perspective Wang et al. (2021); Hellström et al. (2025), $I(\theta)$ tightens the generalization bound by controlling complexity of hypothesis class. The PAC-Bayes bound for a model with parameters $\theta$ is:
$\mathcal{E}_{\text{gen}}(\theta)\leq\mathcal{E}_{\text{emp}}(\theta)+\frac{I(\theta)+\log(1/\delta)}{2n}$
where, $\mathcal{E}_{\text{gen}}(\theta)$ is generalization error and can be defined as: $\mathcal{E}_{\text{gen}}(\theta)=\mathbb{E}_{(x,y)\sim P_{\text{test}}}[\ell(f_{\theta}(x),y)],$ while $\mathcal{E}_{\text{emp}}(\theta)$ is empirical error and is given by: $\mathcal{E}_{\text{emp}}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\ell(f_{\theta}(x_{i}),y_{i}),$ $\theta$ is confidence parameter and $n$ is number of training samples. Minimization of $I(\theta)$, reduce $\mathcal{E}_{\text{gen}}(\theta)$, ensures model robustness to distribution shift.
The second term, CMP given in eq 2 penalizes overconfidence by redistributing log-likelihood to the true class, it ensures that does not assign excessively high probabilities to incorrect classes. To mitigate overconfidence we integrate $\lambda_{2}CMP$ into loss function:$\mathcal{L}_{\text{C3SC}}=\mathcal{L}_{\text{c}}+\lambda_{2}CMP.$ From a Bayesian perspective, CMP act as regularization that encourage the model to distribute probability mass more uniformly across plausible classes. This helps in model calibration by aligning the predicted probabilities with actual correctness likelihood.
From the Minimum Description Length (MDL) principle Hansen and Yu (2001); Kalai and Vempala (2024); Zhao et al. (2024), the CMP reduces the complexity of learned hypothesis by controlling overconfident predictions. The MDL bound for a model with $\theta$ parameters can be defined as:
$\mathcal{E}_{\text{gen}}(\theta)\leq\mathcal{E}_{\text{emp}}(\theta)+\frac{\mathcal{C}(\theta)+\log(1/\delta)}{2n}.$
Where $\mathcal{C}(\theta)$ is hypothesis complexity. By minimizing CMP, model effectively reduces the hypothesis complexity, provides a more compact representation that generalizes better. It helps the model to form robust decision boundaries.
$I(\theta)$ regularization ensures that representations remain robust under domain shifts and CMP ensures that the model does not become overconfident, improving calibration. Since neither penalty degrades the other’s effect, C3SC optimally balances robustness and calibration.
Thus, minimizing the loss function given in eq 3, guarantees a pareto-optimal balance between robustness to covariate shift and also preserving the confidence calibration.
$\mathcal{L}_{\text{$C3SC$}}(x,y;\theta)=\mathcal{L}_{c}+\lambda_{1}I(\theta)+\lambda_{2}\text{CMP}$ (3)
More detail theoretical background about $I(\theta)$ and CMP can be found in in appendix A.1.
3 Experiments and Results
The experimental detail, datasets, and evaluation metrics used in this paper can be found in appendix B. here, we discuss the robustness and confidence calibration results of our method $C3SC$.
C3SC robustness in vision tasks. The results in Table 1 demonstrates the effectiveness of FIM penalty integration into CoOP as compared to baselines. The $\Delta$ row in Table 1 shows improvement across most datasets. The notable improvement in accuracy include Imagenet ($6.8\%$), Flowers102102102102 ($8.0\%$), and UCF101101101101 ($7.3\%$), indicates significant improvement, from action recognition to fine-grained classification. In scene and texture-based datasets $C3SC$ shows robustness performance improvement with SUN397397397397 ($6.0\%$) and DTD ($4.9\%$). For Food101101101101 ($1.7\%$), Caltech101101101101 ($0.2\%$), and OxfordPets ($0.7\%$) improvement in accuracy. $C3SC$ improves in accuracy by $3.2\%$ on average across all datasets.
C3SC calibration in vision tasks. The CMP penalty-based calibration performance results across eleven vision datasets are shown in Table 2. In Table 2, it is shown that CMP integration improves in calibration for most datasets with an average of $5.70\%$ reduction in calibration. The smaller values are better in calibration case. Penalty-based CoOp shows $9.85\%$ calibration improvement at maximum on Food101, mageNet ($8.93\%$), and OxfordPets ($7.19\%$), which demonstrates CMP alignment in model calibration in these settings. On other datasets also it shows better performance such as, UCF101101101101 ($4.55\%$), Caltech101 ($4.94\%$), StanfordCars ($6.51\%$), and FGVCAircraft ($4.94\%$).
In comparison to CLIP, CoOP is surpassed by CLIP but when penalty is added then CoOp shows better calibration. Integration of penalty improves calibration across almost all datasets except Flowers102102102102, where calibration is slightly worsen. The results provides an emprical evidence that penalty integration is reliable technique for reducing overconfidence and improving model reliability in low-shot learning.
Comparing penalty-bassed CoOp to CLIP, penalty-based enhancement provides significant improvement across almost all datasets except StanfordCars and EuroSAT. $C3SC$ shows UCF101101101101 ($14.4\%$), Flowers102102102102 ($13.4\%$), and ImageNet ($12.5\%$) improvement at maximum in comparison to CLIP.
The integration of FIM penalty generally shows improvement accuracy in low-shot vision tasks particular when combined with prompt learning methods.
C3SC robustness in covariate shift. Table 3 shows the robustness performance of $C3SC$ on eight domain generalization datasets under covariate shift. The $\Delta$ of Table 3 shows the penalty-based CoOp outperform CoOp across all datasets, showing significant improvement in robustness to covariate shift. $C3SC$ shows maxmimum robustness performance in Office-Home ($3.5\%$) and VLCS ($3.5\%$) demonstrating FIM integration enhances adaptation across domain adaptation tasks. For all other datasets such as, DomainNet ($1.2\%$), ImageNet-V2 ($1.3\%$), ImageNet-S ($0.9\%$), ImageNet-A ($0.8\%$), and ImageNet-R ($1.6\%$), $C3SC$ shows slight better improvement, suggesting better robustness in challenging real-world distribution shift.
While comparing penalty-based CoOp to CLIP, we observe consistent improvement in accuracy across all domain generalization datasets like, VLCS ($4.6\%$) , Office-Home ($5.2\%$) and PACS ($1.9\%$). $C3SC$ demonstrates better robustness on more challenging variants of ImageNet such as, ImageNet-V2 ($1.3\%$) , ImageNet-S ($2.6\%$), ImageNet-A ($2.7\%$), and ImageNet-R ($2.9\%$).
The FIM penalty indicates consistent performance improvement across covariate shift datasets, reinforcing the FIM penalty role in domain generalization and robustness to covariate shift datasets.
C3SC calibration in covariate shift. Table 4 presents, calibration performance of penalty integration into CoOp for covariate shift vision datasets. It is demonstrated, that C3SC improves in calibration across all datasets with an average of $2.21\%$ at minimum and $5.82\%$ at maximum. C3SC shows calibration improvement on Office-Home ($5.82\%$), PACS ($5.45\%$), VLCS ($5.64\%$), ImageNet-R ($5.45$%) DomainNet ($4.26\%$) and ImageNet-V2 ($3.34\%$), while smaller but consistent improvements are seen in ImageNet-S ($2.62\%$) and ImageNet-A ($2.21\%$), which indicates the CMP effectiveness in confidence alignment on these datasets.
In comparison to CLIP, CLIP shows better calibration as compared to CoOP by reporting lower ECE such as, PACS ($2.18$ vs. $2.02$), VLCS ($2.89$ vs. $3.01$), and ImageNet-R ($3.51$ vs. $3.12$). However, when penalty is integrated into CoOp, the calibration reduces significantly making it more comparable to CLIP.
Our method is effective in confidence alignment as it shows better ECE consistently on almost all datasets. These results favors C3SC calibration reliability under covariate shift, making it more trustworthy in real-world applications.
Table 1: $C3SC$ accuracy results on vision datasets with and without FIM penalty. The $\Delta$ row shows the percentage increase ($\uparrow$) or decrease ($\downarrow$) in accuracy.
ACC Method UCF101 Food101 Caltech101 OxfordPets Flowers102 ImageNet StanfordCars FGVCAircraft SUN397 DTD EuroSAT Avg.
CLIP 69.9 90.1 96.8 91.2 72.1 72.4 63.3 27.2 69.4 53.3 56.5 69.3
CoOp 78.6 97.0 98.6 98.2 79.2 79.5 59.2 25.2 63.0 52.5 53.8 71.2
CoOp + FIM 84.3 98.7 98.8 98.9 85.5 84.9 54.3 27.2 66.8 55.1 49.2 73.5
$\Delta$ 7.3 ↑ 1.7 ↑ 0.2 ↑ 0.7 ↑ 8.0 ↑ 6.8 ↑ 8.2 ↓ 7.9 ↑ 6.0 ↑ 4.9 ↑ 8.6 ↓ 3.2 ↑
Table 2: $C3SC$ ECE results on vision datasets with and without CMP penalty. The $\Delta$ row shows the percentage increase ($\uparrow$) decrease ($\downarrow$) in calibration.
ECE Method UCF101 Food101 Caltech101 OxfordPets Flowers102 ImageNet StanfordCars FGVCAircraft SUN397 DTD EuroSAT Avg.
CLIP 3.24 1.57 6.49 2.25 3.11 1.51 3.74 3.03 1.59 4.53 9.12 3.52
CoOp 3.08 3.35 3.24 3.06 2.96 3.36 3.38 3.24 3.02 3.06 3.08 3.16
CoOp + CMP 2.94 3.02 3.08 2.84 3.16 3.06 3.16 3.08 2.96 2.92 3.06 2.98
$\Delta$ 4.55$\downarrow$ 9.85$\downarrow$ 4.94$\downarrow$ 7.19$\downarrow$ 6.76$\uparrow$ 8.93$\downarrow$ 6.51$\downarrow$ 4.94$\downarrow$ 1.99$\downarrow$ 4.57$\downarrow$ 0.65$\downarrow$ 5.70$\downarrow$
4 Conclusion
This work introduces C3SC a unified framework to address two critical challenges covariate shift and confidence misalignment in low-shot setting vision tasks with foundation model. Integration of Fisher information and CMP, significantly improve both robustness and calibration across various vision and domain generalization datasets. Experimental results demonstrate that C3SC outperforms existing methods in both accuracy and calibration consistently across datasets (upto $3.5\%$) on challenging covariate shift datasets) and calibration (reducing Expected Calibration Error by up to $5.82\%$). These findings highlight C3SC as a promising solution for reliable real-world applications, particularly in low-shot settings where distribution shifts and overconfidence are prevalent.
Author Contributions
If you’d like to, you may include a section for author contributions as is done in many journals. This is optional and at the discretion of the authors.
Acknowledgments
Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
Appendix A Appendix
A.1 Theoretical Background
Proposition A.1.
The Fisher information $I(\theta)$is equal to the Kullback-Leibler (KL) divergence $D_{\text{KL}}$ between the source distribution $p(X)$ and target distribution $q(X)$, providing a measure of covariate shift.
Proof.
Let $p(X;\theta)$ is probability density function parameterized by $\theta$ and $q(X)$ is underlying true distribution. The $D_{\text{KL}}$ between $p(X;\theta)$ and $q(X)$ is given by:
$D_{\text{KL}}(q(X)\parallel p(X;\theta))=\mathbb{E}_{q(X)}\left[\log q(X)-\log p(X;\theta)\right].$
By def. of Fisher information matrix (FIM):
$I(\theta)=-\mathbb{E}_{p(X;\theta)}\left[\frac{\partial^{2}\log P(X;\theta)}{\partial\theta^{2}}\right].$
[…]
$D_{\text{KL}}(q(X)\parallel p(X;\theta))\approx\frac{1}{2}(\theta_{q}-\theta)^{T}I(\theta)(\theta_{q}-\theta).$
where, $\theta_{q}$ is optimal parameter which best fit $q(X)$. By taking second order Taylor expansion of $\log p(X;\theta)$ around $\theta_{q}$:
$\log p(X;\theta)\approx\log p(X;\theta_{q})+(\theta-\theta_{q})^{T}\nabla_{\theta}\log p(X;\theta_{q})+\frac{1}{2}(\theta-\theta_{q})^{T}H(\theta-\theta_{q}),$
where $H=\nabla^{2}_{\theta}\log p(X;\theta)$ is the Hessian.
By taking expectation w.r.t to $q(X)$, the first term vanishes as Fisher information expectation is defined over second order derivative. The $D_{\text{KL}}$ simplifies to:
$D_{\text{KL}}(q(X)\parallel p(X;\theta))\approx\frac{1}{2}(\theta_{q}-\theta)^{T}I(\theta)(\theta_{q}-\theta).$
Thus, Fisher information provide a measure of divergence between $q(X)$ and $p(X;\theta)$, capturing covariate shift. ∎
Proposition A.2.
The Confidence Misalignment Penalty (CMP) provides the measure for degree of overconfidence in incorrect predictions and also ensures alignment between predicted and true confidence score. CMP ensures that it is always a valid confidence $0\leq\text{CMP}\leq 1$, $\text{CMP}\approx 1$ when model is overconfidence, and $\text{CMP}\approx 0$ when the model is well-calibrated.
Proof.
Let CMP be defined as:
$\text{CMP}=\frac{P(x,y^{\prime})}{\sum_{y_{j}\neq y^{\prime}\,:\,P(x,y_{j})>P(x,y^{\prime})}P(x,y_{j})}$
where, $P(x,y^{\prime})$ is probability of predicted class $y^{\prime}$. The denominator considers only $y_{j}$ where $P(x,y_{j})>P(x,y^{\prime})$, which means that CMP is small when all other classes have higher confidence than true class $y^{\prime}$.
If the model is highly confident with incorrect prediction, then $P(x,y^{\prime})$ is high, it can be expressed mathematically as:
$\lim_{P(x,y^{\prime})\to 1}\text{CMP}=1$
In case of well-calibration, the confidence mass is distributed across all other classes, reducing $P(x,y^{\prime})$, ensuring $\text{CMP}\to 0$:
$\lim_{P(x,y^{\prime})\to 0}\text{CMP}=0$
Since all probabilities sum to 1, CMP is ramins always within range $[0,1]$, making it a valid calibration measure.
$0\leq\text{CMP}\leq 1$
Thus, CMP acts as confidence misalignment penalty, discouraging overconfident yet incorrect predictions. ∎
Proposition A.3.
Integration of CMP and $I(\theta)$ into loss function ensures covariate shift correction and confidence calibration as well.
Proof.
The final modified CLIP loss given by:
$\mathcal{L}_{\text{$C3SC$}}(x,y;\theta)=\mathcal{L}_{c}+\lambda_{1}I(\theta)+\lambda_{2}\text{CMP}$
Where, As per theorem A.1, $I(\theta)$ addresses covariate shift, while proposition A.2 ensures confidence calibration.
In gradient update, $I(\theta)$ adjust the weight updates based on local curvature, prevents divergence due to covariate shift:
$\Delta\theta\propto I(\theta)^{-1}\nabla\mathcal{L}(\theta)$
where $I(\theta)^{-1}$ is scaling gradient to account local curvature while $\mathcal{L}(\theta)$ is the loss function.
CMP reduces the magnitude of updates to address the high confidence:
$\nabla\text{CMP}\cdot\nabla\mathcal{L}(\theta)\approx 0\quad\text{when confidence is high.}$.
If $\lambda_{2}>0$, then CMP ensure probability mass is distributed across classes rather than overly concentrating on incorrect prediction:
$\lambda_{2}>0\implies\sum_{y\neq y^{\prime}}P(x,y)$
This implies that when $\lambda_{2}>0$ then probability mass is distributed among classes. Tuning $\lambda_{1},\lambda_{2}$ properly prevents overconfidence without penalising well-aligned distributions.
Thus, minimizing $\mathcal{L}_{\text{$C3SC$}}$ ensures well-calibrated robust prediction under covariate shift. ∎
Appendix B Experiments
Datasets. To check the effectiveness of $C3SC$, we used 19 vision dataset including natural distribution datasets, domain adaptation datasets and vision classification datasets used for evaluation by low-shot prompt learning methods. These datasets are: Caltech101 Fei-Fei et al. [2004], Imagenet Deng et al. [2009], EuroSAT Helber et al. [2019], StanfordCars Krause et al. [2013], FGVCAircraft Maji et al. [2013], OxfordPets Parkhi et al. [2012], SUN397 Xiao et al. [2010], Food101 Bossard et al. [2014], Flowers102 Nilsback and Zisserman [2008], DTD Cimpoi et al. [2014], UCF101 Soomro [2012], PACS Li et al. [2017], VLCS Fang et al. [2013], Office-Home Venkataraman et al. [2016] and five domain adaptation datasets such as Imagent-1k and its variants like ImageNet A (Art) Hendrycks et al. [2021a], ImageNet-V2 (V2) Recht et al. [2019], ImageNet R (Real) Hendrycks et al. [2021b] and ImageNet S (Sketch) Wang et al. [2019].
Baselines. We consider CLIP and prompt learning method CoOpZhou et al. [2022] without any penalty integration as baseline in comparison to our method.
Evaluation metrics. We used accuracy (Acc) as evaluation metric for robustness of $C3SC$ under covariate shift and vision datasets used by prompt learning approaches, while Expected Calibration Error (ECE) Guo et al. [2017] is used with same number of datasets for evaluating the calibration performance of $C3SC$.
Implementation details. To check the robustness and calibration performance of $C3SC$, we conduct extensive experiments on various datasets. We used pre-trained CLIP Radford et al. [2021] with ViT-B/16 Dosovitskiy [2020] as backbone in prompt learning method. In covariate shift experiment we used CLIP with ViT-B/16 and ResNet-50 backbone He et al. [2016] for the last four variants of ImageNet. For first four datasets we follow the same protocol of Li et al. [2017], where model is trained on another domain and evaluated on another test domain.
Table 3: $C3SC$ accuracy results on covariate shift vision datasets with and without FIM penalty. The $\Delta$ row shows the percentage increase/decrease in accuracy. $\uparrow$ shows improvement in accuracy while the $\downarrow$ shows decrease in performance.
ACC Method PACS Office-Home VLCS DomainNet ImageNet-V2 ImageNet-S ImageNet-A ImageNet-R
CLIP 96.1 80.4 81.4 54.1 60.8 46.2 47.8 73.9
CoOp 96.5 82.1 82.5 58.8 64.2 47.9 49.7 75.2
CoOp+FIM 98.0 85.6 86.0 60.0 65.5 48.8 50.5 76.8
$\Delta$ 1.5 $\uparrow$ 3.5 $\uparrow$ 3.5 $\uparrow$ 1.2 $\uparrow$ 1.3 $\uparrow$ 0.9 $\uparrow$ 0.8 $\uparrow$ 1.6 $\uparrow$
Table 4: $C3SC$ ECE results on covariate shift vision datasets with and without CMP penalty. The $\Delta$ row shows the percentage increase/decrease in calibration. $\downarrow$ shows improvement in calibration while $\uparrow$ the shows calibration performance decrease.
ECE Method PACS Office-Home VLCS DomainNet ImageNet-V2 ImageNet-S ImageNet-A ImageNet-R
CLIP 2.18 2.77 2.89 3.69 2.44 4.88 8.34 3.51
CoOp 2.02 2.92 3.01 3.29 4.19 8.40 15.34 3.12
CoOp+CMP 1.91 2.75 2.84 3.15 4.05 8.18 15.00 2.95
$\Delta$ 5.45 $\downarrow$ 5.82 $\downarrow$ 5.64 $\downarrow$ 4.26 $\downarrow$ 3.34 $\downarrow$ 2.62 $\downarrow$ 2.21 $\downarrow$ 5.45 $\downarrow$
Table 5: Accuracy three datasets for tuning $\boldsymbol{\lambda_{1}}$ within range ($0.0-1.0$) while keeping $\boldsymbol{\lambda_{2}}$ value fixed $0.1$.
$\boldsymbol{\lambda_{1}}$ Datasets
Flowers102102102102 Food101101101101 UCF101101101101 DTD
0.0 83.5 96.2 83.5 53.1
0.2 81.3 94.8 83.1 53.5
0.4 85.5 98.7 84.3 55.1
0.6 84.7 90.8 80.4 52.2
0.8 82.4 89.1 82.9 52.2
1.0 81.8 87.0 80.9 51.5